OpenStudy (anonymous):

If $ 2500 is invested in an account that pays interest compounded continuously, how long will it take to grow to $ 5000 at 3%? Round to the nearest tenth.

OpenStudy (anonymous):

For comtinuously compounded interest we can use \[ P(t) = P _{o}e ^{rt},\]where\[P _{0}\]is the initial deposit, r is the annual percentage interest rate, and t is in years.

OpenStudy (anonymous):

A=Pe^(rt) A=2P so the Ps cancel 2=e^9.03t) ln both sides ln2=.03t divide by .03 1=23.1

OpenStudy (anonymous):

That's what I got.